OpenStudy (anonymous):

Which of the following is a solution of 4x2 = - 9x - 4?

pooja195 (pooja195):

Do the directions specify what we have to solve by? Quadratic formula or Factoring?

pooja195 (pooja195):

Step one: We need to get the equation into \[\huge~\rm~ax^2+bx+c=0 \] form. What do we do to get the equation into this form?

OpenStudy (anonymous):

Quadratic formula

OpenStudy (anonymous):

would i minus the 4x^2 on both sides?

pooja195 (pooja195):

yes! :)

OpenStudy (anonymous):

ok now what ?

pooja195 (pooja195):

\[\huge~\rm-4x^2- 9x - 4=0\]

pooja195 (pooja195):

list your A B C values now :)

OpenStudy (anonymous):

a.9 plus or minus the square root of 17 divided by 8. b.negative 9 plus or minus the square root of 145 divided by 8. c.9 plus or minus the square root of 145 divided by 8. d.negative 9 plus or minus the square root of 17 divided by 8

pooja195 (pooja195):

abc values...

OpenStudy (anonymous):

i dont understand:/

pooja195 (pooja195):

ax^2+bx+c=0 -4x^2-9x-4=0 name the abc values now

OpenStudy (anonymous):

a=-4 b=-9 c=-4? like that

pooja195 (pooja195):

Yes :) \[\huge~\rm~~x=\frac{ -(-9)~\pm \sqrt{(-9)^2-4(-4)(-4)} }{ 2(-4) } \]

pooja195 (pooja195):

Solve \[\huge~\rm~~x=\frac{ 9~\pm \sqrt{(-9)^2-4(-4)(-4)} }{ 2(-4) } \]

OpenStudy (anonymous):

\[\frac{ 81-64? }/-8

OpenStudy (anonymous):

is this right?

pooja195 (pooja195):

Not quite look at this: http://prntscr.com/7mvafb

OpenStudy (anonymous):

that makes a lot more since lol thank you so much. i got tones more but ill use your method lol